# Knowledge Graph Retrieval-Augmented Generation for LLM-based Recommendation

### v1

@@ 4 Experiment
In this section, we evaluate the effectiveness of our proposed framework through comprehensive experiments. First, we present the experimental settings, including details about the datasets, compared baselines, evaluation metrics, and the parameter configurations. Then, we report the main experimental results, highlighting the performance of the proposed framework compared with various baseline methods. Finally, we analyze the contributions of individual model components and the impact of parameters used in our framework. We also assess the generalizability of K-RagRec in the zero-shot setting.
@@ 4.1.1 Datasets
Table 1: Basic statistics of three datasets and the KG. ”Items in KG” indicates the number of items that appeared in both the KG and the dataset.
@@ 4.1.2 Baselines
In the realm of LLM-based recommendation research, our work pioneers the investigation of retrieving knowledge from KGs to enhance the recommendation capabilities of LLMs. Therefore, to evaluate the effectiveness, we compare our proposed framework with a series of meticulously crafted KG RAG-enhanced LLM recommendation baselines. We first include two typical inference-only methods Retrieve-Rewrite-Answer Wu et al. (2023b) (KG-Text) and KAPING Baek et al. (2023), where the former retrieves sub-graphs and textualizes them, and the latter retrieves triples. We exclude some knowledge reasoning path-based approaches Luo et al. (2023), as it is difficult to retrieve faithful knowledge reasoning paths solely from the user’s interaction items. Next, we compare K-RagRec with various prompt-tuning approaches augmented by retrieval, including Prompt Tuning with KG-Text (PT w/ KG-Text), GraphToken Perozzi et al. (2024) with retrieval (GraphToken w/ RAG) as well as G-retriever He et al. (2024). Additionally, we evaluate our method against Lora Fine-tuning with retrieval (Lora w/ KG-Text) Hu et al. (2021).
@@ 4.1.3 Evaluation Metrics
To evaluate the effectiveness of our K-RagRec framework, we employ two widely used evaluation metrics: Accuracy (ACC), and Recall@$k$  He et al. (2020). We present results for $k$ equal to 3, and 5. Inspired by recent studies Zhang et al. (2024); Hou et al. (2022), we adopt the leave-one-out strategy for evaluation. Specifically, for each user, we select the last item that the user interacted with as the target item and the 10 interaction items prior to the target item as the historical interactions. Then, we leverage LLM to predict the user’s preferred item from a pool of 20 candidate items ($M=20$), which contains one target item with nineteen randomly sampled items. For trained models (including prompt tuning and fine-tuning), we compute Recall@$k$ by extracting the probability assigned to each item and evaluating the model’s ability to rank the target item within the top-$k$ predictions. In addition, we also conduct comparison experiments with 10 candidate items ($M=10$) as shown in Appendix A.2.
@@ 4.1.4 Parameter Settings
We implement the framework on the basis of PyTorch and conduct the experiments on 2 NVIDIA A6000-48GB GPUs. We adopt the SentenceBert to encode entities, relations, and query attributes. We use the 3 layers Graph Transformer as the $\text{GNN}^{\text{Indexing}}$ and $\text{GNN}^{\text{Encoding}}$ for MovieLens-1M and 4 layers for MovieLens-20M and Amazon Book. The layer dimension is set to 1024, and the head number is set to 4. The adaptive retrieval policy threshold $p$ is set to 50%. For each item that needs to be retrieved, we retrieve the top-3 most similar sub-graphs. The re-ranking knowledge sub-graph number $N$ is set to 5. More experiment details are shown in Appendix A.1. We also present several prompt examples in Appendix A.3.
Models Methods MovieLens-1M MovieLens-20M Amazon Book ACC R@3 R@5 ACC R@3 R@5 ACC R@3 R@5 LLama-2 Inference-only KG-Text Wu et al. (2023b) 0.076 - - 0.052 - - 0.058 - - KAPING Baek et al. (2023) 0.079 - - 0.069 - - 0.063 - - Frozen LLM w/ PT PT w/ KG-Text 0.078 0.191 0.308 0.051 0.152 0.250 0.074 0.165 0.245 GraphToken w/ RAG Perozzi et al. (2024) 0.268 0.421 0.466 0.186 0.433 0.576 0.326 0.515 0.624 G-retriever He et al. (2024) 0.274 0.532 0.650 0.342 0.619 0.739 0.275 0.487 0.612 K-RagRec 0.435 0.725 0.831 0.600 0.850 0.913 0.508 0.690 0.780 Improvement 58.6% 33.0% 27.8% 75.4% 37.3% 23.5% 55.8% 34.0% 25.0 % Fine-tuning Lora w/ KG-Text 0.402 0.718 0.833 0.609 0.848 0.905 0.446 0.648 0.758 Lora w/ K-RagRec 0.466 0.770 0.863 0.637 0.872 0.927 0.516 0.720 0.799 Improvement 15.9% 7.2% 3.6% 4.5% 2.7% 2.4% 15.7% 11.1% 5.4% LLama-3 Inference-only KG-Text Wu et al. (2023b) 0.095 - - 0.060 - - 0.054 - - KAPING Baek et al. (2023) 0.084 - - 0.069 - - 0.062 - - Frozen LLM w/ PT PT w/ KG-Text 0.134 0.294 0.433 0.094 0.205 0.296 0.083 0.207 0.314 GraphToken w/ RAG Perozzi et al. (2024) 0.355 0.622 0.737 0.473 0.719 0.805 0.428 0.567 0.661 G-retriever He et al. (2024) 0.352 0.632 0.746 0.502 0.736 0.796 0.417 0.584 0.682 K-RagRec 0.472 0.704 0.765 0.634 0.779 0.818 0.514 0.662 0.723 Improvement 32.9% 11.4% 2.5% 26.3% 5.8% 1.6% 20.0% 13.4% 6.0% Fine-tuning Lora w/ KG-Text 0.449 0.694 0.750 0.648 0.757 0.790 0.490 0.638 0.698 Lora w/ K-RagRec 0.498 0.712 0.771 0.674 0.786 0.817 0.546 0.672 0.733 Improvement 10.9% 2.6% 2.8% 4.0% 3.8% 3.4% 11.4% 5.3% 5.0% QWEN2 Inference-only KG-Text Wu et al. (2023b) 0.160 - - 0.174 - - 0.194 - - KAPING Baek et al. (2023) 0.196 - - 0.208 - - 0.220 - - Frozen LLM w/ PT PT w/ KG-Text 0.190 0.371 0.499 0.259 0.397 0.494 0.303 0.451 0.553 GraphToken w/ RAG Perozzi et al. (2024) 0.259 0.487 0.608 0.370 0.550 0.632 0.365 0.568 0.658 G-retriever He et al. (2024) 0.304 0.551 0.644 0.389 0.606 0.685 0.355 0.552 0.658 K-RagRec 0.416 0.712 0.829 0.586 0.842 0.904 0.502 0.686 0.767 Improvement 36.8% 29.2% 28.4% 50.6% 38.9% 32.0% 37.5% 20.8% 16.6 % Fine-tuning Lora w/ KG-Text 0.400 0.701 0.815 0.601 0.842 0.906 0.478 0.667 0.751 Lora w/ K-RagRec 0.466 0.763 0.860 0.631 0.868 0.928 0.510 0.704 0.780 Improvement 16.5% 8.8% 5.5% 5.0% 3.1 % 2.4% 6.7% 5.5% 3.9%
@@ 4.2 Overall Performance Comparison
We compare the recommendation performance of K-RagRec with various baselines on three open-source backbone LLMs: LLama-2-7b Touvron et al. (2023), LLama-3-8b Dubey et al. (2024), and QWEN2-7b Yang et al. (2024). We present the results in Table 2. From the comparison, we have the following main observations:
@@
  • Compared to the LoRA fine-tuning with the naive RAG approach, K-RagRec with prompt tuning can achieve close to or even better performance in most settings. Notably, K-RagRec can achieve the best performance when fine-tuned with LoRA fine-tuning.
4.3 Ablation Study
To evaluate the impact of each component in our proposed framework, we conduct the ablation study to compare the K-RagRec with four ablated variants on MovieLens and Amazon Book datasets, using LLama-2-7B as the backbone LLM. Specifically, we compare the framework with four ablated variants: K-RagRec (-Indexing), K-RagRec (-Adaptive), K-RagRec (-Re-ranking), and K-RagRec (-Encoding). (1) K-RagRec (-Indexing) eliminates the $\text{GNN}^{\text{Indexing}}$ and stores semantic information of PLM in the knowledge vector database. For the retrieved nodes, we extract their second-order sub-graphs as the retrieved knowledge sub-graphs. (2) K-RagRec (-Adaptive) does not apply the adaptive retrieval policy and retrieves all items from the user’s historical interactions. (3) (-Re-ranking) removes the re-ranking module and inputs the knowledge sub-graphs directly. (4) K-RagRec (-Encoding) removes the $\text{GNN}^{\text{Encoding}}$ and replaces it with a trainable soft prompt. The retrieved knowledge sub-graphs will be added to the prompt as triples (e.g., {Moonraker, film writer film, Christopher Wood (writer)}).
The results are illustrated in Figure 3. Observing the experiment results, we find that eliminating any component of the framework leads to a decrease in the overall performance of the recommendations, demonstrating the effectiveness of each module. Secondly, removing the GNN Encoder leads to a 37% decrease and a 45.9% decrease in the accuracy of the model on MovieLens and Amazon Book datasets, respectively, highlighting the significance of employing GNN to encode the structure of knowledge sub-graphs.
@@ 4.3 Ablation Study
4.4 Generalization Study
Models Methods MovieLens-20M Amazon Book
ACC R@3 R@5 ACC R@3 R@5
LLama2 K-RagRec 0.539 0.740 0.795 0.390 0.581 0.671
Lora w/K-RagRec 0.539 0.783 0.863 0.405 0.580 0.796
LLama3 K-RagRec 0.597 0.797 0.839 0.428 0.628 0.706
Lora w/K-RagRec 0.611 0.775 0.814 0.424 0.622 0.732
QWEN K-RagRec 0.507 0.769 0.861 0.418 0.612 0.687
Lora w/K-RagRec 0.545 0.814 0.897 0.441 0.623 0.706
@@ 4.4 Generalization Study
GNN Types ACC R@3 R@5
GCN Kipf & Welling (2016) 0.397 0.704 0.809
GAT Velickovic et al. (2017) 0.420 0.693 0.804
Graph Transformer Shi et al. (2020) 0.429 0.711 0.779
GraphSAGE Hamilton et al. (2017) 0.418 0.699 0.823
(a) Accuracy
(b) Recall@5
(c) Inference Time
Figure 4: Effect of adaptive retrieval policy threshold $p$ on MovieLens-1M and LLama-2-7b across metrics Accuracy, Recall@5 and inference time (seconds) for K-RagRec.
@@ 4.5 Efficiency Evaluation
In this sub-section, we evaluate the inference efficiency of our proposed K-RagRec framework compared with baselines on the MovieLens-1M dataset and LLama-2-7b. We record the time cost for one inference utilizing two NVIDIA A6000-48G GPUs. The time cost for a single inference is reported in Table 4. By observing the experimental results, we notice that various KG RAG approaches significantly increase the inference time, which is due to the large scale of the KG. In contrast, K-RagRec achieves the best computational efficiency compared to various KG RAG methods and is only about 0.1s slower than direct inference without retrieval. These findings highlight the efficiency of K-RagRec and validate the effectiveness of our adaptive retrieval policy.
4.6 Study of Four GNN Encoders
To further understand the generality of our proposed approach, we conduct a comparative study of four variants applying different GNN encoders. Specifically, we compare GCN (Kipf & Welling, 2016), GAT (Velickovic et al., 2017), Graph Transformer (Shi et al., 2020), and GraphSAGE (Hamilton et al., 2017) as four K-RagRec variants of GNN encoder. Specifically, Graph Convolutional Network (GCN) (Kipf & Welling, 2016) first introduces convolutional operations to graph-structured data. By aggregating features from neighboring nodes, GCN facilitates the learning of rich node representations. GraphSAGE (Hamilton et al., 2017) learns an aggregation function that samples and combines features from a node’s local neighborhood in an inductive setting, enabling the effective use of new nodes. Graph Attention Network (GAT) (Velickovic et al., 2017) further incorporates attention mechanisms, allowing the model to dynamically assign varying attention to neighboring nodes, thereby enhancing the focus on the most relevant information. Inspired by the success of the transformer, the Graph Transformer (Shi et al., 2020) adapts transformer architectures to graph data, enhancing the modeling of graphs, particularly textual graphs.
@@ 4.7 Parameter Analysis
In this section, we evaluate the impact of four main hyper parameters of K-RagRec, namely adaptive retrieval policy threshold $p$, retrieved knowledge sub-graph numbers $K$, re-ranking knowledge sub-graph numbers $N$, and GNN layer numbers.
(a) Amazon Book: Accuracy
(b) Amazon Book: Recall@3
(c) Amazon Book: Recall@5
(a) Amazon Book: Accuracy
(b) Amazon Book: Recall@3
(c) Amazon Book: Recall@5
1) Impact of adaptive retrieval policy threshold $p$: To understand how the adaptive retrieval policy threshold $p$ affects K-RagRec, we conduct experiments on MovieLens-1M and LLama-2-7b across two metrics. Results are shown in Figure 4. When threshold $p$ is set to a small value, where only a few items are augmented, leading to insufficient retrieval and reduced recommendation accuracy. On the other hand, as observed in Figure 4 (c), the inference time increases almost linearly with threshold $p$. Therefore, selecting an appropriate threshold $p$ is crucial to balance the performance and inference time.
@@
Table 6: Comparison of the impact of different GNN layers on the Amazon Book dataset and LLama-2-7b across three metrics. ACC and R@k denote Accuracy and Recall@k, respectively.
GNN Layers ACC R@3 R@5
3 layers 0.496 0.653 0.736
4 layers 0.506 0.690 0.780
5 layers 0.498 0.656 0.729
3) Impact of GNN Layer Numbers: In this part, we evaluate the impact of the number of GNN layers on model performance. We vary the numbers of the GNN layer numbers in the range of {3, 4, 5} and test on the Amazon Book dataset and LLama-2-7b across three metrics. As observed in Table 6, the model performance first improves and then decreases as the number of GNN layers increases, and the model achieves the best results across the three metrics when setting the number of GNN layers is set to four. Therefore, a smaller number of GNN layers may not have sufficient depth to capture the intricate relationships and dependencies in the graph, leading to sub-optimal performance. On the other hand, too many layers can result in indistinguishable node representations. Thus, selecting the optimal number of GNN layers is crucial for effective model training.
@@ 5 Conclusion
In this paper, we propose a novel framework K-RagRec to augment the recommendation capability of LLMs by retrieving reliable and up-to-date knowledge from KGs. Specifically, we first introduce a GNN and PLM to perform semantic indexing of the KGs, enabling both coarse and fine-grained retrieval for KGs. To further improve retrieval efficiency and effectiveness, we introduce an adaptive retrieval policy that determines whether an item needs to be retrieved based on its popularity. Notably, K-RagRec performs more expressive graph encoding of the retrieved knowledge sub-graphs, facilitating the LLM to effectively leverage the structure information and avoid long context input. Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework.

### v2

@@ 4 Experiment
In this section, we evaluate the effectiveness of our proposed framework through comprehensive experiments. First, we present the experimental settings, including details about the datasets, compared baselines, evaluation metrics, and the parameter configurations. Then, we report the main experimental results, highlighting the performance of the proposed framework compared with various baseline methods. Finally, we analyze the contributions of individual model components and the impact of parameters used in our framework. We also assess the generalizability of K-RagRec in the zero-shot setting. We present the generalizability study in Appendix A.7.
@@ 4.1.2 Baselines
In the realm of LLM-based recommendation research, our work pioneers the investigation of retrieving knowledge from KGs to enhance the recommendation capabilities of LLMs. Therefore, to evaluate the effectiveness, we compare our proposed framework with a series of meticulously crafted KG RAG-enhanced LLM recommendation baselines. We first include two typical inference-only methods Retrieve-Rewrite-Answer (Wu et al., 2023b) (KG-Text) and KAPING (Baek et al., 2023), where the former retrieves sub-graphs and textualizes them, and the latter retrieves triples. We exclude some knowledge reasoning path-based approaches (Luo et al., 2023), as it is difficult to retrieve faithful knowledge reasoning paths solely from the user’s interaction items. Next, we compare K-RagRec with various prompt-tuning approaches augmented by retrieval, including Prompt Tuning with KG-Text (PT w/ KG-Text), GraphToken (Perozzi et al., 2024) with retrieval (GraphToken w/ RAG) as well as G-retriever (He et al., 2024). Additionally, we evaluate our method against Lora Fine-tuning with retrieval (Lora w/ KG-Text) (Hu et al., 2021).
@@ 4.1.3 Evaluation Metrics
To evaluate the effectiveness of our K-RagRec framework, we employ two widely used evaluation metrics: Accuracy (ACC), and Recall@$k$  (He et al., 2020). We present results for $k$ equal to 3, and 5. Inspired by recent studies (Zhang et al., 2024; Hou et al., 2022), we adopt the leave-one-out strategy for evaluation. Specifically, for each user, we select the last item that the user interacted with as the target item and the 10 interaction items prior to the target item as the historical interactions. Then, we leverage LLM to predict the user’s preferred item from a pool of 20 candidate items ($M=20$), which contains one target item with nineteen randomly sampled items. For trained models (including prompt tuning and fine-tuning), we compute Recall@$k$ by extracting the probability assigned to each item and evaluating the model’s ability to rank the target item within the top-$k$ predictions. In addition, we also conduct comparison experiments with 10 candidate items ($M=10$) as shown in Appendix A.5.
Models Methods MovieLens-1M MovieLens-20M Amazon Book ACC R@3 R@5 ACC R@3 R@5 ACC R@3 R@5 LLama-2 Inference-only KG-Text (Wu et al., 2023b) 0.076 - - 0.052 - - 0.058 - - KAPING (Baek et al., 2023) 0.079 - - 0.069 - - 0.063 - - Frozen LLM w/ PT PT w/ KG-Text 0.078 0.191 0.308 0.051 0.152 0.250 0.074 0.165 0.245 GraphToken w/ RAG (Perozzi et al., 2024) 0.268 0.421 0.466 0.186 0.433 0.576 0.326 0.515 0.624 G-retriever (He et al., 2024) 0.274 0.532 0.650 0.342 0.619 0.739 0.275 0.487 0.612 K-RagRec 0.435 0.725 0.831 0.600 0.850 0.913 0.508 0.690 0.780 Improvement 58.6% 33.0% 27.8% 75.4% 37.3% 23.5% 55.8% 34.0% 25.0 % Fine-tuning Lora w/ KG-Text 0.402 0.718 0.833 0.609 0.848 0.905 0.446 0.648 0.758 Lora w/ K-RagRec 0.466 0.770 0.863 0.637 0.872 0.927 0.516 0.720 0.799 Improvement 15.9% 7.2% 3.6% 4.5% 2.7% 2.4% 15.7% 11.1% 5.4% LLama-3 Inference-only KG-Text (Wu et al., 2023b) 0.095 - - 0.060 - - 0.054 - - KAPING (Baek et al., 2023) 0.084 - - 0.069 - - 0.062 - - Frozen LLM w/ PT PT w/ KG-Text 0.134 0.294 0.433 0.094 0.205 0.296 0.083 0.207 0.314 GraphToken w/ RAG (Perozzi et al., 2024) 0.355 0.622 0.737 0.473 0.719 0.805 0.428 0.567 0.661 G-retriever (He et al., 2024) 0.352 0.632 0.746 0.502 0.736 0.796 0.417 0.584 0.682 K-RagRec 0.472 0.704 0.765 0.634 0.779 0.818 0.514 0.662 0.723 Improvement 32.9% 11.4% 2.5% 26.3% 5.8% 1.6% 20.0% 13.4% 6.0% Fine-tuning Lora w/ KG-Text 0.449 0.694 0.750 0.648 0.757 0.790 0.490 0.638 0.698 Lora w/ K-RagRec 0.498 0.712 0.771 0.674 0.786 0.817 0.546 0.672 0.733 Improvement 10.9% 2.6% 2.8% 4.0% 3.8% 3.4% 11.4% 5.3% 5.0% QWEN2 Inference-only KG-Text (Wu et al., 2023b) 0.160 - - 0.174 - - 0.194 - - KAPING (Baek et al., 2023) 0.196 - - 0.208 - - 0.220 - - Frozen LLM w/ PT PT w/ KG-Text 0.190 0.371 0.499 0.259 0.397 0.494 0.303 0.451 0.553 GraphToken w/ RAG (Perozzi et al., 2024) 0.259 0.487 0.608 0.370 0.550 0.632 0.365 0.568 0.658 G-retriever (He et al., 2024) 0.304 0.551 0.644 0.389 0.606 0.685 0.355 0.552 0.658 K-RagRec 0.416 0.712 0.829 0.586 0.842 0.904 0.502 0.686 0.767 Improvement 36.8% 29.2% 28.4% 50.6% 38.9% 32.0% 37.5% 20.8% 16.6 % Fine-tuning Lora w/ KG-Text 0.400 0.701 0.815 0.601 0.842 0.906 0.478 0.667 0.751 Lora w/ K-RagRec 0.466 0.763 0.860 0.631 0.868 0.928 0.510 0.704 0.780 Improvement 16.5% 8.8% 5.5% 5.0% 3.1 % 2.4% 6.7% 5.5% 3.9%
@@ 4.1.4 Parameter Settings
We implement the framework on the basis of PyTorch and conduct the experiments on 2 NVIDIA A6000-48GB GPUs. We adopt the SentenceBert to encode entities, relations, and query attributes. We use the 3 layers Graph Transformer as the $\text{GNN}^{\text{Indexing}}$ and $\text{GNN}^{\text{Encoding}}$ for MovieLens-1M and 4 layers for MovieLens-20M and Amazon Book. The layer dimension is set to 1024, and the head number is set to 4. The popularity selective retrieval policy threshold $p$ is set to 50%. For each item that needs to be retrieved, we retrieve the top-3 most similar sub-graphs. The re-ranking knowledge sub-graph number $N$ is set to 5. More experiment details are shown in Appendix A.1. We also present several prompt examples in Appendix A.12.
@@ 4.2 Overall Performance Comparison
We compare the recommendation performance of K-RagRec with various baselines on three open-source backbone LLMs: LLama-2-7b (Touvron et al., 2023), LLama-3-8b (Dubey et al., 2024), and QWEN2-7b (Yang et al., 2024). We present the results in Table 1. From the comparison, we have the following main observations:
@@
  • Compared to the LoRA fine-tuning with the naive RAG approach, K-RagRec with prompt tuning achieves close to or better performance in most settings. Notably, K-RagRec achieves the best performance when fine-tuned with LoRA.
@@
4.3 Ablation Study
To evaluate the impact of each component in our proposed framework, we conduct the ablation study to compare the K-RagRec with four ablated variants on MovieLens and Amazon Book datasets, using LLama-2-7B as the backbone LLM. Details of each ablation variant are provided in Appendix A.6. The results are illustrated in Figure 3. Observing the experiment results, we find that eliminating any component of the framework leads to a decrease in the overall performance of the recommendations, demonstrating the effectiveness of each module. Secondly, removing the GNN Encoder leads to a 37% decrease and a 45.9% decrease in the accuracy of the model on MovieLens and Amazon Book datasets, respectively, highlighting the significance of employing GNN to encode the structure of knowledge sub-graphs. Refer to Appendix A.6 for more details.
@@ 4.4 Efficiency Evaluation
In this sub-section, we evaluate the inference efficiency of our proposed K-RagRec framework compared with baselines on the MovieLens-1M dataset and LLama-2-7b. We record the time cost for one inference utilizing two NVIDIA A6000-48G GPUs. The time cost for a single inference is reported in Table 2. By observing the experimental results, we notice that various KG RAG approaches significantly increase the inference time, which is due to the large scale of the KG. In contrast, K-RagRec achieves the best computational efficiency compared to various KG RAG methods and is only about 0.1s slower than direct inference without retrieval. These findings highlight the efficiency of K-RagRec and validate the effectiveness of our popularity selective retrieval policy.
@@ 4.5 Parameter Analysis
In this section, we evaluate the impact of three main hyper parameters of K-RagRec, namely popularity selective retrieval policy threshold $p$, retrieved knowledge sub-graph numbers $K$, and re-ranking knowledge sub-graph numbers $N$. In addition, we analyze the impact of various GNN encoder variants and different GNN layer numbers for the proposed framework in Appendix A.10 and  A.11.
1) Impact of popularity selective retrieval policy threshold $p$: To understand how the popularity selective retrieval policy threshold $p$ affects K-RagRec, we conduct experiments on MovieLens-1M and LLama-2-7b across two metrics. Results are shown in Figure 4. As the threshold $p$ increases, the recommendation performance initially improves and then decreases. When the threshold $p$ is set to a small value, only a few items are augmented. This leads to insufficient retrieval and poor recommendation accuracy. When $p$ is set to a larger value, more items are retrieved for augmentation. However, due to the re-ranking sub-graph numbers $N$ being fixed, some retrieved cold-start item knowledge sub-graphs are discarded or ranked at the back of the list, resulting in sub-optimal recommendation performance. On the other hand, as observed in Figure 4 (c), the inference time increases almost linearly with threshold $p$. Therefore, selecting an appropriate threshold $p$ is crucial to balance the performance and inference time.
@@ 5 Conclusion
In this paper, we propose a novel framework K-RagRec to augment the recommendation capability of LLMs by retrieving reliable and up-to-date knowledge from KGs. Specifically, we first introduce a GNN and PLM to perform semantic indexing of KGs, enabling both coarse and fine-grained retrieval for KGs. To further improve retrieval efficiency, we introduce a popularity selective retrieval policy that determines whether an item needs to be retrieved based on its popularity. Notably, K-RagRec performs more expressive graph encoding of the retrieved knowledge sub-graphs, facilitating the LLM to effectively leverage the structure information and avoid long context input. Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework.